Rogala Foods Inc. was formed in 2015 with the merger of Grouch Mayer and Tashamo Corporation. The company reported the following rounded amounts for the year ended January 3, 2016 (all amounts in millions): 


DebitsCredits
Accounts Receivable$1,100 
Allowance for Doubtful Accounts
$ 28
Sales (assume all on credit)
18, 100


 Required:

 1. Assume Rogala uses 1/2 of 1 percent of sales to estimate its Bad Debt Expense for the year. Prepare the adjusting journal entry required for the year, assuming no Bad Debt Expense has been recorded yet.

 2. Assume instead Rogala uses the aging of accounts receivable method and estimates that $71 of its Accounts Receivable will be uncollectible. Prepare the adjusting journal entry required at January 3, 2016, for recording Bad Debt Expense.

 3. Assume instead that Rogala's uses the aging of accounts receivable method and estimates that $71 of its Accounts Receivable will be uncollectible. Prepare the year-end adjusting journal entry for recording Bad Debt Expense assuming Rogala's unadjusted balance in Allowance for Doubtful Accounts at January 3, 2016 was a debit balance of $32.

 4. If one of Rogala's customers declared bankruptcy, what journal entry would be used to write off its $20 balance?
Answer #1
1
Debit Credit
January 03,2016 Bad Debts expense 91 =18100*1%*1/2
      Allowance for Doubtful Debts 91
2
Debit Credit
January 03,2016 Bad Debts expense 43 =71-28
      Allowance for Doubtful Debts 43
3
Debit Credit
January 03,2016 Bad Debts expense 103 =71+32
      Allowance for Doubtful Debts 103
4
Debit Credit
1 Allowance for Doubtful Debts 20
       Accounts Receivable 20

> you my nigga bro

Zion's Remedy Wed, Dec 1, 2021 8:36 AM